A firm starting a business is trying to decide whether to purchase or lease a phone system. Leasing will cost $6000 per year, and the firm must pay an additional $0.08 per call. Purchasing a phone system will cost $15000 per year, plus $1000 per year for a service contract. With either  system, a long-distance provider will be needed. At what number of calls will purchasing a phone system becomes worthwhile?
Answer #1

Let number of calls be x

6000+ 0.08(x)= 15000+1000

x= (16000-6000)/0.08

x= 125000

If calls are more than 125000, then purchasing a phone system is worthwhile.